# Human-Like Robot Impedance Regulation Skill Learning from Human-Human Demonstrations

### v1

@@ IV-B1 Task Definition
The first task is human-robot collaborative transportation, which consists of teaching the CURI to manipulate and transport an object by following the human partner while dealing with the position and impedance constraints (as shown in Fig. 7). Specifically, they will collaboratively lift, translate, and lower the object to the target pose under leader-follower mode. It is worth noting that the start and the end poses of the object are not fixed during demonstrations. We propose to teach the cobot follower to perform consistent compliance behavior as the human leader (shown in Fig. 2).
We use the CURI as the cobot which was introduced in Section III-D as shown in Fig. 3. The robotic hand of qb SoftHand2 Research was applied as the end effector. For the software development, a whole-body Cartesian impedance control interface was implemented based on section III-D in ROS. By giving the commanded impedance matrices and the reference trajectory with an update frequency of 100 $Hz$, the controller outputted the commanded torques of each joint for the robotic arms and calculated joint velocities for torso joints to generate expected motions with a frequency of 1000 $Hz$. Note that the gravity compensation was redesigned due to the unique installations of the dual arms.
@@ IV-B2 Trajectory and Impedance Generation from Demonstration
The demonstration data of the collaborative transportation task were recorded by the Optitrack system, EMG sensors, and force sensors with the corresponding software. The demonstrators’ motions were obtained from the Optitrack markers, which were fixed on the human joints at a frequency of 120 $Hz$ (shown in Fig. 7). The data were used for offline learning to generate a proper representation of the cobot motions. Besides, the demonstrators were wearing two EMG sensors on the upper body (same positions as the calibration process) to record the muscle activities. The recording and processing procedures of EMG signals were the same as the calibration part. These data were used for human endpoint stiffness estimation during the demonstration. The reference trajectories were then calculated based on the human endpoint motion, stiffness, and interactive force by (8).
As shown in Fig. 8, we utilized three calculated reference trajectories of human demonstrations in GMM. The result of the representation with regard to the PoE was illustrated for the object transportation task. Furthermore, by adding the task parameter and extending to TP-GMM, the motion reproduction was achieved with new initial and final poses of the reference trajectory and the generated reference trajectory was given. Besides, the corresponding stiffness variables along the trajectory were calculated based on the LQT method. Finally, the LSTM model was built to learn the regulation policy between the two demonstrators (shown in Fig. 5).
Figure 9: The snapshots of human-robot collaborative object transportation with our proposed method and a constant impedance method. (We declare that the individual in this figure is author and he gave permission for the use of his image.)
Figure 10: The comparison of cobot movement and interactive force on CURI right end effector between the constant impedance method (orange solid line) and our proposed method (blue solid line) in three directions. The results of two different trials are given.
The collection and preprocessing of the multimodal demonstration data such as IMU and EMG signals, and the implementation of the TP-GMM and LQT methods depend on the Rofunc package, a full-process python package for robot learning from demonstration published by our lab [40].
@@ IV-B3 Experimental Results
In real world human-robot experiments shown in Fig. 9, the human-like cobot, CURI, was equipped with an Optitrack marker behind its torso to obtain the poses of the end effector in the world coordinates. These data were transferred into a rosbag format through data streaming. Hence, the movements of CURI were transformed into the cobot coordinates by mapping between the marker and the end effector. The execution of whole-body motion (arm and torso in this paper) was based on the reference trajectories and the corresponding impedance profiles of the end effector learned offline from demonstration via TP-LQT method. The human partner was wearing the EMG sensors during the experiments to monitor the muscle co-contraction, which was used to calculate the regulatory parameters of the cobot based on the regulation policy.
The snapshots of the human-robot collaborative object carrying task with our proposed method were given in the upper part of Fig. 9 with both side and top views. The cobot and the human partner successfully lifted the object from the initial position, translated the object, and lowered the object to the target position. The movement of the human-robot collaboration in these three stages was similar to the human-human demonstration. We also performed a comparative experiment with the same reference trajectory while the impedance of the end effector was set as constant ($K^{d}$ = 300 $N/m$ and $D_{d}$ = 34.64 $N/{m^{2}}$) in all three directions. Note that it was a common way to verify the effectiveness of the proposed method [41]. The snapshots of the experimental results were illustrated in the lower part of Fig. 9. We found that the object was pulled too much to the cobot side during the translating stage, which made the object deviate from the demonstration range (indicated by the red solid circles in Fig. 9).
Figure 11: The snapshots of human-robot confrontational Tai Chi pushing hands with both side and top views. II.A: Tai Chi unimanual flat circle pushing hand, II.B: Tai Chi bimanual vertical circle pushing hands. (We declare that the individual in this figure is author and he gave permission for the use of his image.)
We have recorded the interactive force from the cobot side by the force sensor to further verify the advantage of our approach during the experiments. The end effector movement in Cartesian space was also calculated based on the forward kinematics of the CURI. The results were plotted in Fig. 10, which include two trials with different start and end poses. We can find out that the movements generated by the constant impedance method were closer to the cobot side, which caused the issue mentioned above. When comparing the interaction force of the cobot end effector during the collaboration, the trends were smoother by using our proposed approach. The average interactive forces in three directions were also calculated and shown in Table II. The average force on each axe was similar between the two methods in trial 1 while the average interaction force of our proposed method was smaller than the constant one (especially on the $Z$-axes) in trial 2. This phenomenon shows the advantage of the proposed method to some extent.
TABLE II: average interactive forces comparison
Trial 1 Trial 2
constant proposed ($N$) constant proposed ($N$)
X 2.08 1.78 2.75 1.26
Y 3.19 2.54 2.46 2.81
Z 3.80 2.74 7.63 3.29
Figure 12: The comparison of average interactive forces on CURI end effectors between a constant impedance control method with orange line and our proposed method with blue line in three directions. II.A: Tai Chi unimanual flat circle pushing hand, II.B: Tai Chi bimanual vertical circle pushing hands.
IV-C Tai Chi Pushing Hands Task
@@ IV-C1 Task Definition
The second task is Tai Chi pushing hands, which is widely known as a popular Chinese kong-fu because of its health benefits, especially for elders. As the basis of the pushing hands, the fixed step involves pushing motions in flat circles, vertical circles, and ‘8’-shaped circles [42]. In this paper, unimanual pushing of flat circles and bimanual pushing of vertical circles are selected for human-robot collaboration shown in Fig. 7.
The physical behaviors of the two individuals being confronted were discovered to be mutually adaptive during Tai Chi pushing hands (see Fig. 7). More specifically, when showing pushing motions, one individual has a stiff behavior to interacting by hands while the other performs quite compliant on the first half of the motion path. Conversely, the roles of these two individuals have switched on the return. Hence, the cobot has to generate regulatory compliance behavior compared to the human partner in each stage of Tai Chi, which will be learned from the impedance regulation module.
@@ IV-C2 Trajectory and Impedance Generation from Demonstration
The data collection setup of confrontational Tai Chi pushing hands was the same as the collaborative transportation task (shown in Fig. 7). Two sub-tasks including the unimanual flat circle and bimanual vertical circle pushing hands were performed by two human demonstrators. Since we only calibrated one upper arm of each demonstrator, the parameters of the other arm were set to the same values as the calibrated one for simplification in the bimanual task.
The motion representation and generation results are shown in Fig. 8. Besides, impedance regulation policies of both Tai Chi tasks were learned by the LSTM networks and one example of training and testing predictions was given in Fig. 5. The trained models were further applied to get the regulatory factors for the cobot to tune compliance during HRC.
@@ IV-C3 Experimental Results
In the real-world experiments (as illustrated in Fig. 11), the setup of the cobot and the human partner was almost the same as the first one except two robotic arms were applied in the second Tai Chi task.
@@
The snapshots of two human-robot confrontational pushing hands with unimanual flat circle pushing and bimanual vertical circle pushing were shown in Fig. 11. In each motion cycle, CURI utilized the learned regulation policy and achieved the task within 10 $s$. During the unimanual arm pushing task, CURI initially exerted high impedance to push the human towards one side and subsequently reduced the pushing force in the latter half of the movement. In the bimanual pushing task, CURI dual arms adjusted their actions based on the co-contraction of human muscles, transitioning between different stages. For the two Tai Chi pushing hands tasks, the stiffness and damping variables of one or two end effectors in Cartesian space were modified according to regulation policies. Specifically, during the unimanual flat circle pushing, CURI right arm maintained high stiffness and damping in the initial phase and decreased the impedance in the latter half. Similarly, in the bimanual pushing tasks, the impedance variations mirrored those of the right arm with stage transitions, while exhibiting opposite behaviors in the left arm.
Our proposed method was evaluated against a constant impedance control method through four repetitions of these human-robot collaboration tasks. Comparative analysis of the average interactive forces between the end effector and the human hand in three directions relative to the impedance frame of the robotic arm is presented in Fig. 12. In both tasks, the trends in average interactive forces were similar. However, the amplitudes of these forces were greater when utilizing our proposed method, compared to the smaller forces observed with the constant impedance control method, particularly in the directions of movement. This experimental outcome demonstrates that our method enables the cobot to exhibit more appropriate physical behaviors during human-robot collaboration, thereby facilitating more effective interactions between humans and cobots.

### v2

@@ IV-B1 Task Definition
The first scenario is human-robot collaborative transportation, where the cobots are expected to help the human partners by manipulating and transporting the objects while dealing with the position and impedance constraints (as shown in Fig. 5.I). Specifically, human and robot will collaboratively lift, translate, and lower the object to the target pose under leader-follower mode. It is worth noting that the start and the end poses of the object are not fixed during collaborations. In this task, we propose to teach the cobot follower to perform consistent compliance behavior as the human leader (shown in Fig. 2).
One arm and the torso of CURI were used for collaborative transportation tasks. A robotic hand from qb SoftHand2 Research was applied as the end effector for grasping. A whole-body Cartesian impedance control interface was implemented based on section III-D in ROS. By giving the commanded impedance matrices and the reference trajectory with an update frequency of 100 $Hz$, the controller outputted the commanded torques of each joint for the robotic arms and calculated joint velocities for torso joints to generate expected motions with a frequency of 1000 $Hz$. Note that the gravity compensation was redesigned due to the unique installations of the arms.
Figure 8: The snapshots of human-robot collaborative tasks by HI-ImpRSL method. I: object transportation, II: Tai Chi pushing hands, and III: sawing.
@@ IV-B2 Trajectory and Impedance Generation from Demonstration
The demonstration data of the collaborative transportation were recorded by the OptiTrack system, EMG sensors, and force sensors with the corresponding software. The demonstrators’ motions were obtained from the OptiTrack markers, which were fixed on the human joints (shown in Fig. 5). The data were used for offline learning to generate a proper representation of the cobot motions. Besides, the demonstrators were wearing two EMG sensors on the upper body (same positions as the calibration process) to record the muscle activities. The recording and processing procedures of EMG signals were the same as the calibration part. These data were used for human endpoint stiffness estimation. The reference trajectories were then calculated based on the human endpoint motion, stiffness, and interactive force by (8).
As shown in Fig. 6.I, we utilized three calculated reference trajectories of human demonstrations in GMM. The result of the representation with regard to the PoE was illustrated for the object transportation task. Furthermore, by adding the task parameter and extending to TP-GMM, the motion reproduction was achieved with new initial and final poses of the reference trajectory and the generated reference trajectory was given. Besides, the corresponding stiffness variables along the trajectory were calculated based on the LQT method.
The LSTM model was presented in the upper part of Fig. 7, which compares the predicted muscle synergistic contributions with the ground truth values of the collaborator during an object transportation task. The demonstration data from Subject 1 (S1, collaborator) and Subject 5 (S5, partner) were used as the training set for constructing the model, while additional subject combinations (S1 and S3, S1 and S4) were utilized as testing sets. The accuracy of the LSTM network predictions was evaluated using the Root Mean Squared Error (RMSE) and the correlation coefficient. The results indicate that the prediction trends align well with the ground truth, with performance metrics for the testing sets (S1&S3: RMSE 17.67, correlation 0.7307; S1&S4: RMSE 15.99, correlation 0.6383) slightly lower than those of the training set (RMSE 11.84, correlation 0.7928). The observed performance decline in the testing sets can be attributed to the inherent variability of EMG signals across different individuals. Given these differences, such a reduction in prediction accuracy is acceptable as long as the model successfully captures the overall trends, which is crucial for generalizing to diverse interaction scenarios.
@@ IV-B3 Experimental Results
In real world human-robot experiments shown in Fig. 8, the human-like cobot, CURI, was equipped with an OptiTrack marker behind its torso to obtain the poses of the end effector in the world coordinates. These data were transferred into a ROS bag format through data streaming. Hence, the movements of CURI were transformed into the cobot coordinates by mapping between the marker and the end effector. The execution of whole-body motion was based on the reference trajectories and the corresponding impedance profiles of the end effector learned offline from demonstration via TP-LQT method. The human partner was wearing the EMG sensors during the experiments to monitor the muscle co-contraction, which was used to calculate the regulatory parameters of the cobot based on the LSTM-based regulation policy. To control expectancy and effort biases, each subject completed 25 randomized trials comprising five impedance control methods (five trials per method), with method identity masked and trial order fully randomized across different subjects.
The snapshots of the human-robot collaborative object transportation task with HI-ImpRSL method were given in Fig. 8 with two subjects. The cobot and the human partners successfully lifted the object from the initial position, translated the object, and lowered the object to different target positions. The movement of the human-robot collaboration in these three stages was similar to the human-human demonstration. We also performed four comparative experiments with the same reference trajectory while the impedance variations of the end effector were different. The FIC method set $K_{d}$ = 300 $N/m$ and $D_{d}$ = 34.64 $N\cdot s/{m}$ in all three directions. The EMG-based method integrates biosignals to adapt impedance parameters based on task-specific rules, which was defined as a positive linear correlation in the leader-follower mode. The NN-based VIC method applied the interaction force-based stiffness estimation and used the variable impedance skill framework, which was learned from demonstrations based on a neural network. The HI-ImpRSL w/o LSTM method directly used the impedance profiles which learned from human demonstrations without the LSTM module.
TABLE II: Interactive Forces Statistics Comparison of Task I (Mean $\pm$ SD)
Method Force Mean (N) Smoothness ($10^{-3}$ $N/s^{2}$)
X Y Z X Y Z
FIC 1.66 $\pm$ 0.27 1.10 $\pm$ 0.10 5.82 $\pm$ 0.82 2.19 $\pm$ 0.11 2.23 $\pm$ 0.13 2.30 $\pm$ 0.25
EMG-based VIC 1.54 $\pm$ 0.29 1.11 $\pm$ 0.20 5.15 $\pm$ 0.92 2.17 $\pm$ 0.17 2.24 $\pm$ 0.16 2.62 $\pm$ 0.27
NN-based VIC 1.24 $\pm$ 0.32 0.99 $\pm$ 0.18 5.05 $\pm$ 0.57 2.01 $\pm$ 0.23 1.98 $\pm$ 0.17 2.27 $\pm$ 0.20
HI-ImpRSL w/o LSTM 1.34 $\pm$ 0.24 1.01 $\pm$ 0.12 5.53 $\pm$ 0.84 2.25 $\pm$ 0.30 2.22 $\pm$ 0.31 2.36 $\pm$ 0.17
HI-ImpRSL 1.18 $\pm$ 0.27 0.92 $\pm$ 0.12 5.27 $\pm$ 0.27 1.94 $\pm$ 0.20 1.86 $\pm$ 0.23 2.06 $\pm$ 0.17
*Force Mean: the time-averaged absolute force magnitude; Smoothness: weighted RMSD of force jerk (second derivative);
*Values are aggregated over subjects S2–S5 as mean $\pm$ standard deviation.
We have recorded the interaction force from the cobot-side force/torque sensor to further verify the advantage of our method during the experiments. Unless otherwise specified, the reported “mean interaction force” is computed as the time-averaged absolute force magnitude. The results between CURI and subjects were shown in Table II.
The subject-averaged results were given among S2-S5, while HI-ImpRSL delivers the best overall performance in both interaction force magnitude and smoothness. For force magnitude, HI-ImpRSL attains the lowest values in the $X$ and $Y$ directions ($1.18\pm 0.27N$ and $0.92\pm 0.12N$, respectively), outperforming FIC, EMG-based VIC, and the ablated variant HI-ImpRSL w/o LSTM ($1.34\pm 0.24N$ and $1.01\pm 0.12N$, respectively). In the $Z$ direction, NN-based VIC achieves the lowest mean force ($5.05\pm 0.57N$), while HI-ImpRSL remains competitive ($5.27\pm 0.27N$).
In terms of smoothness, quantified as the weighted RMSD of the second derivative of the interaction forces, HI-ImpRSL consistently achieves the smallest values across all axes ($1.94\pm 0.20\cdot 10^{-3}N/s^{2}$, $1.86\pm 0.23\cdot 10^{-3}N/s^{2}$, and $2.06\pm 0.17\cdot 10^{-3}N/s^{2}$ in $X$, $Y$, and $Z$ direction). These results indicate reduced high-frequency fluctuations and improved interaction stability compared with FIC, EMG-based VIC, NN-based VIC ($2.01\pm 0.23\cdot 10^{-3}N/s^{2}$, $1.98\pm 0.17\cdot 10^{-3}N/s^{2}$, and $2.27\pm 0.20\cdot 10^{-3}N/s^{2}$ in $X$, $Y$, and $Z$ direction), and HI-ImpRSL w/o LSTM. Overall, the aggregated statistics suggest that HI-ImpRSL most effectively suppresses interaction disturbances while maintaining the smoothest force profiles among the compared methods.
IV-C Tai Chi Pushing Hands
@@ IV-C1 Task Definition
The second task is Tai Chi pushing hands, which is widely known as a popular Chinese kung fu because of its health benefits, especially for elders. As the basis of the pushing hands, the fixed step involves pushing motions in flat circles, vertical circles, and ‘8’-shaped circles [wang2016biomechanical]. We select unimanual pushing of flat circles and bimanual pushing of vertical circles for collaboration (Fig. 5).
Tai Chi pushing hands inherently features mutual adaptation, with partners alternating between stiff and compliance across the motion cycle. One leads with a stiffer push while the other yields compliantly, and the roles reverse on the return. Framed in rehabilitation terms, this alternating pattern aligns with passive–active training, where therapist-like guidance is paired with patient-like yielding. Hence, the cobot has to generate regulatory compliance behavior compared to the human partner in each stage of Tai Chi, which will be learned from the impedance regulation module.
TABLE III: Interactive Forces Statistics Comparison of Task II and III (Mean $\pm$ SD)
Task Method Force Mean (N) Smoothness ($10^{-3}$ N/s2)
X Y Z X Y Z
Tai Chi Uni FIC 20.47 $\pm$ 7.40 6.00 $\pm$ 2.10 6.77 $\pm$ 3.12 5.40 $\pm$ 1.02 4.87 $\pm$ 0.82 5.52 $\pm$ 1.15
EMG-based VIC 15.46 $\pm$ 6.60 5.78 $\pm$ 2.04 4.97 $\pm$ 1.40 5.15 $\pm$ 1.11 4.76 $\pm$ 0.59 3.95 $\pm$ 0.62
NN-based VIC 22.54 $\pm$ 9.34 4.98 $\pm$ 1.87 4.25 $\pm$ 1.07 4.98 $\pm$ 0.96 4.14 $\pm$ 0.52 2.88 $\pm$ 0.61
HI-ImpRSL w/o LSTM 18.76 $\pm$ 6.47 5.51 $\pm$ 2.14 4.14 $\pm$ 0.83 5.11 $\pm$ 0.89 5.12 $\pm$ 1.62 4.56 $\pm$ 1.04
HI-ImpRSL 19.97 $\pm$ 8.34 4.66 $\pm$ 1.53 3.98 $\pm$ 1.24 5.26 $\pm$ 1.28 3.61 $\pm$ 0.36 2.45 $\pm$ 0.16
Tai Chi Bi: Right FIC 4.27 $\pm$ 0.98 4.82 $\pm$ 2.14 17.48 $\pm$ 3.35 3.93 $\pm$ 0.40 4.46 $\pm$ 0.27 4.97 $\pm$ 1.72
EMG-based VIC 4.43 $\pm$ 1.66 4.62 $\pm$ 2.36 16.65 $\pm$ 4.05 3.82 $\pm$ 0.49 4.34 $\pm$ 0.47 5.45 $\pm$ 1.79
HI-ImpRSL w/o LSTM 3.08 $\pm$ 0.73 4.20 $\pm$ 2.00 16.84 $\pm$ 3.10 3.46 $\pm$ 0.17 4.66 $\pm$ 1.48 4.98 $\pm$ 1.41
HI-ImpRSL 2.70 $\pm$ 0.56 4.16 $\pm$ 1.80 16.09 $\pm$ 3.14 2.97 $\pm$ 0.26 3.68 $\pm$ 0.37 4.89 $\pm$ 2.41
Tai Chi Bi: Left FIC 2.95 $\pm$ 1.01 4.55 $\pm$ 1.90 13.07 $\pm$ 5.78 4.04 $\pm$ 0.46 4.65 $\pm$ 1.10 6.19 $\pm$ 2.50
EMG-based VIC 2.96 $\pm$ 0.36 3.95 $\pm$ 0.43 12.35 $\pm$ 2.57 4.10 $\pm$ 0.35 4.49 $\pm$ 1.24 5.36 $\pm$ 2.05
[…]
HI-ImpRSL 2.28 $\pm$ 0.69 3.08 $\pm$ 0.91 13.32 $\pm$ 4.78 2.65 $\pm$ 0.34 3.33 $\pm$ 0.53 4.82 $\pm$ 1.92
Sawing FIC 5.51 $\pm$ 6.71 3.21 $\pm$ 1.37 6.80 $\pm$ 3.86 8.45 $\pm$ 2.18 5.07 $\pm$ 0.97 10.26 $\pm$ 6.19
EMG-based VIC 5.74 $\pm$ 6.30 2.95 $\pm$ 1.67 7.13 $\pm$ 3.65 6.72 $\pm$ 3.82 4.33 $\pm$ 0.89 9.13 $\pm$ 5.64
[…]
HI-ImpRSL 7.71 $\pm$ 8.13 1.62 $\pm$ 0.96 9.18 $\pm$ 2.98 6.32 $\pm$ 4.38 3.12 $\pm$ 0.34 7.47 $\pm$ 3.52
@@ IV-C2 Trajectory and Impedance Generation from Demonstration
The data collection setup of confrontational Tai Chi pushing hands was the same as the collaborative transportation task. Two sub-tasks including the unimanual flat circle (Fig. 5.II) and bimanual vertical circle (Fig. 5.III) pushing hands were performed by two human demonstrators. Since we only calibrated one upper arm of each demonstrator, the parameters of the other arm were set to the same values as the calibrated one for simplification in the bimanual task.
The motion representation and generation results of the unimanual Tai Chi pushing tasks are shown in Fig. 6.II. Besides, impedance regulation policies of both Tai Chi tasks were learned by the LSTM networks and one example of training and testing predictions was given in Fig. 7.II. The trained models were further applied to get the regulatory factors for the cobot to tune compliance.
@@ IV-C3 Experimental Results
In the real-world experiments (as illustrated in Fig. 8), the setup of the cobot and the human partner was almost the same as the first task, except that two robotic arms were applied in the second Tai Chi task. Four impedance regulation methods for the cobot were performed with every subject in random order, which included 5 trials for each method and a total of 20 trials for each subject.
@@
The snapshots of human-robot confrontational pushing hands with unimanual flat circle pushing and bimanual vertical circle pushing were shown in Fig. 8. In each motion cycle, CURI utilized the learned regulation policy and achieved the task. For instance, during the unimanual arm pushing task, CURI initially exerted high impedance to push the human towards one side and subsequently reduced the pushing force in the latter half of the movement. Specifically, CURI right arm maintained high stiffness and damping in the initial phase and decreased the impedance in the latter half.
The statistical analysis of interactive forces with four subjects is given in Table III, which illustrates the average performance of S2-S5 in the form of MEAN $\pm$ SD. Our proposed HI-ImpRSL exhibits the most favorable overall performance in both interaction force magnitude and smoothness across the collaborative Tai Chi tasks. For the unimanual task, where the motion primarily lies in the $XY$ plane and the $Z$-axis force best reflects disturbance rejection, HI-ImpRSL attains the lowest $Z$-direction mean force ($3.98\pm 1.24N$), outperforming FIC ($6.77\pm 3.12N$), EMG-based VIC ($4.97\pm 1.40N$), NN-based VIC ($4.25\pm 1.07N$), and HI-ImpRSL w/o LSTM ($4.14\pm 0.83N$). It also achieves the smallest $Y$-direction force ($4.66\pm 1.53N$), indicating improved suppression of out-of-plane disturbances and higher cooperativity.
For bimanual tasks, where the trajectory is mainly in the $YZ$ plane and the force on the $X$-axis is most indicative, HI-ImpRSL achieves the lowest mean forces in the $X$-direction on both arms. On the right arm, it records $2.70\pm 0.56N$, outperforming FIC ($4.27\pm 0.98N$), EMG-based VIC ($4.43\pm 1.66N$), and HI-ImpRSL w/o LSTM ($3.08\pm 0.73N$). On the left arm, it achieves $2.28\pm 0.69N$, lower than FIC ($2.95\pm 1.01N$), EMG-based VIC ($2.96\pm 0.36N$), and HI-ImpRSL w/o LSTM ($3.17\pm 1.24N$). These reductions corroborate HI-ImpRSL’s effectiveness in minimizing interaction forces and enhancing bilateral coordination.
Smoothness, quantified as the weighted RMSD of the second derivatives of the force signals, further underscores the stability of HI-ImpRSL. In the unimanual task, it achieves the lowest values along $Y$ and $Z$ ($3.61\pm 0.36\cdot 10^{-3}N/s^{2}$ and $2.45\pm 0.16\cdot 10^{-3}N/s^{2}$, respectively), outperforming all baselines. Besides, NN-based VIC attains the smallest value on $X$ ($4.98\pm 0.96\cdot 10^{-3}N/s^{2}$) while HI-ImpRSL remains competitive ($5.26\pm 1.28\cdot 10^{-3}N/s^{2}$). For bimanual tasks, HI-ImpRSL consistently yields the lowest smoothness metrics on all axes on both the right arm ($2.97\pm 0.26,3.68\pm 0.37,4.89\pm 2.41\cdot 10^{-3}N/s^{2}$) and the left arm ($2.65\pm 0.34,3.33\pm 0.53,4.82\pm 1.92\cdot 10^{-3}N/s^{2}$). Overall, the aggregated statistics indicate that HI-ImpRSL most effectively suppresses interaction disturbances while maintaining the smoothest force profiles across the evaluated Tai Chi tasks.
IV-D Collaborative Sawing
We further introduce a collaborative sawing task to evaluate mutual adaptation under tool-environment dynamics. A human and the cobot jointly operate a hand saw to cut a wooden block (as shown in Fig. 8). We adopt a hybrid position-force scheme that decouples planar stroke generation from normal-force regulation. In the $XY$ plane, a reference trajectory is learned via TP-LQT from demonstrations and adapted online to the current frame. The cobot’s in-plane stiffness and damping are modulated by the partner’s muscle activation, using processed EMG envelopes to adjust impedance gains for phase-aware compliance and stability under varying friction. Along the $Z$ axis, a force controller maintains a constant normal force (10NN) to ensure cutting engagement while compensating surface irregularities. Interaction forces are measured at the end effector, and EMG provides a proxy of human effort for impedance regulation.
Quantitatively, HI-ImpRSL attains the smallest off-axis interaction forces in the $Y$ direction ($1.62\pm 0.96N$) compared to the other three methods. Regarding the smoothness, our method consistently achieves smoother interaction ($3.12\pm 0.34\cdot 10^{-3}N/s^{2}$). Along $Z$ direction, where the controller regulates a 10 $N$ normal force, HI-ImpRSL exhibits $9.18\pm 2.98N$ (underlined as closest to the 10 $N$ reference), demonstrating superior force-tracking accuracy compared with FIC ($6.80\pm 3.86N$), EMG-based VIC $(7.13\pm 3.65N$), and the ablation ($7.67\pm 4.08N$). These improvements confirm that HI-ImpRSL simultaneously enhances planar compliance and normal-force fidelity in the collaborative sawing task.
Figure 9: The muscle activation comparison between HI-ImpRSL and demonstration, including collaborative transportation (top) and Tai Chi unimanual pushing (bottom) performed by S3 and S4.
IV-E Discussion
To compare the collaboration dynamics of the leader-follower and mutual adaptation modes, we analyzed the smoothness and variations of the interaction forces recorded during human-robot collaboration. In the transportation task, interaction forces were smoother and exhibited lower variations in the main direction of movement. In contrast, the Tai Chi pushing hands task showed greater variability due to its mutual adaptation mode, highlighting the task’s bidirectional compliance regulation and dynamic role switching. These results demonstrate the contrasting interaction dynamics between the two types of tasks, with the leader-follower emphasizing stable compliance and the mutual adaptation requiring more adaptive and variable behaviors, validating the framework’s adaptability across different collaboration modes.
We further compared the muscle activation trends of S3 and S4 during the demonstrations and HRC with our HI-ImpRSL framework across the collaborative transportation task and the Tai Chi unimanual pushing task (Fig. 9). The results revealed that the overall trends in muscle activation were highly similar, further validating that our proposed method can effectively generate human-like compliance behaviors. However, we observed a notable difference in the initial phase of the Tai Chi task for S3, where the muscle activation levels were less aligned. This discrepancy is attributed to the tendency to exert force during the preparatory phase when collaborating with a human partner, whereas with the robotic system, force application occurred only after the motion had begun. Aside from this, the subsequent muscle activation trends were consistent.
It is worth noting that beyond Tai Chi pushing hands, our proposed HI-ImpRSL can be extended to rehabilitation or medical robotics by using human-human therapeutic demonstrations to derive task-specific trajectories and impedance profiles. The LSTM-based regulation can adapt assistance or resistance to patient effort with safety-aware compliance, indicating broad clinical applicability.